OpenStudy (anonymous):

Find the slope of the graph of the function at F(x) =4/x^2 when x=5

OpenStudy (anonymous):

f(5)=4/5^2=5/25

OpenStudy (anonymous):

Oh, that's all? Thanks so much!

OpenStudy (anonymous):

But I need to find f'(5), not just f(5)

OpenStudy (anonymous):

F(x) =4/x^2 y =4/x^2 to find the inverse you need to change the x and y then solve for y x = 4/y^2 y^2*x=4 y^2=4/x sqrt(y^2)=sqrt(4)/sqrt(x) y = 2/sqrt(x) = 2sqrt(x)/x

OpenStudy (anonymous):

2*sqrt(5)/5

OpenStudy (anonymous):

does that look right?

OpenStudy (anonymous):

errr, I don't think so, sorry : / These are my answer choices: A. -8/5 B. 8/125 C. 8/5 D. -8/125

OpenStudy (anonymous):

by f' u mean derivative or inverse

OpenStudy (anonymous):

f=4/x^2 = 4x^(-2) f'=-2*4*x^-3= -8/x^3 f'(5) =-8/5^3=-8/125

OpenStudy (anonymous):

Ohhh, that makes so much sense. Thanks!!

OpenStudy (anonymous):

For some reason, I can never remember how to rewrite exponents that are in fractions. Or whenever a root is involved.

OpenStudy (anonymous):

How would you do 3x^2+5x-6/x^2 when x=-5? My answer choices end up with a fraction with 3137 as the numerator somehow